# Towards Immersive Mixed Reality Street Play

### v1

@@ 3.3. System Implementation
To address the inaccessibility of expensive MRHMDs described in Section 2.5.1, we use HoloKit [^13] https://holokit.io/products/holokit-x, visited on October 25, 2024, an affordable (\$129) smartphone-based open-source stereoscopic optical see-through mixed reality headset. This cost-effective solution enables multiplayer mixed reality experiences ”in the wild,” allowing a large number of passersby to try the gameplay.
We implemented the game probes using Unity 6 LTS [^14] https://unity.com/releases/unity-6, visited on October 25, 2024 and the HoloKit SDK [^15] https://github.com/holokit/holokit-unity-sdk, visited on October 25, 2024 based on ARKit [^16] https://developer.apple.com/augmented-reality/arkit/, visited on October 25, 2024. We deployed the game as an iOS application on iPhone 12 or later models, which are inserted into the HoloKit. The headset displays stereoscopic renderings on the iPhone’s screen, allowing users to see stereoscopic 3D images overlaid on reality when viewed through HoloKit.
We achieved interactions with augmented content through an Apple Watch. The watch’s motion sensors capture wrist movements for spell casting, enabling remote attacking gameplay without physical contact between players. This approach cleverly sidesteps the physical contact issue by allowing players to cast magic spells for intangible interaction with opponents. MOFA’s system features various spell classes, each with a basic attack spell and a utility spell. Players select one magic class per game round. While basic attack spells have minor variations in speed and size, utility spells differ significantly, adding depth to the gameplay.
[…]
@@ 4. Method
[…]
To comprehensively investigate the dynamics and societal acceptance of immersive mixed reality street play, we employed a mixed-methods approach. We collected data through onsite behavioral observations and interviews using the ”research in-the-wild” method (Benford et al., 2013a), and supplemented this with online reviews using desktop research techniques.
[…]
To investigate IMRSP, we conducted an in-the-wild study (Benford et al., 2013a) combining direct gameplay video observations with immediate post-play contextual inquiry interviews. To gain a broader perspective, we shared video recordings of the street play on an online forum. We also conducted desktop research to gather and analyze online reviews and comments about the experience. This dual approach allowed us to capture both firsthand player experiences and wider public perceptions. Our study was approved by the authors’ institutional ethics review, and all participants gave informed consent prior to participation.
@@ 4.1. In-the-Wild Field Study
[…]
Figure 3. Five scenarios of volunteered participation and interviews. S1: Campus; S2: Expos; S3: Urban parks; S4: Public buildings; S5: Natural wildness.
[…]
@@ 4.1.1. Procedure
[…]
In each of these pop-up shows, two researchers initially conducted the game in a space of at least 3x3 meters in a public space, ensuring minimal disruption to pedestrian flow. Upon attracting passersby who stopped and started to observe, researchers offered them opportunities to try the game. Interested participants then consent to participate in the game and the data collection before using the headsets. Participants could play for their desired duration, typically completing two rounds, with the first round serving as a tutorial. During the game, researchers observed player behaviors and the sessions were video recorded. Where time and availability of the participants allowed, researchers invited them to join a semi-structured interview guided by the following questions: The interviews were conducted either individually or in small groups of 2-3 players who had participated together. On average, each participant played the game for 4.08 minutes, and the interview lasted for 2.75 minutes.
[…]
In order to collect ecologically valid MR street play data, we used an opportunistic recruitment strategy. This intercept recruitment approach is common in-the-wild research as expected, it yielded relatively short engagements due to the nature of interrupting people during their normal routine. Nevertheless, it provided a broad range of people who encountered this MR street play in those public spaces, enhancing the diversity of our sample. A detailed procedure of an example street play is described below.
@@
    Configure the Play Area: In each pop-up or in-booth show, two or three researchers conducted the game in a space of at least 3×3 meters in a public area, ensuring minimal disruption to pedestrian flow. A “Video Recording In Progress” signage is displayed. For the expo scenarios (S2), the game was demonstrated both at the booth and through pop-up installations in lobbies or near entrances during breaks, following the in-the-wild deployment approach.
@@
    Engage Passersby: Pedestrians were drawn to the ongoing gameplay, introduction videos, and spectator footage displayed on iPads and portable screens. As they approached to watch and interact with researchers, they were offered opportunities to try the game themselves.
@@
    Obtain Informed Consent: Interested participants signed consent forms on an iPad or on a hard copy.
@@
[…]
@@
    • When another participant became available, they were often invited to join in “The Duel”, which requires 2 players and encourages competitive interaction. Participants who arrived together—such as friends or colleagues—were directly introduced to The Duel to leverage their existing social rapport.
    • For participants with more time and physical availability, and when the play area permitted extended movement, researchers suggested “The Dragon”, a longer (about 5 minutes) and more physically demanding cooperative game mode.
@@
    Facilitate Gameplay Sessions: The first 90 seconds, or the first round, was a tutorial session. After that, when the participants learned how to attack and doge intuitively in MR, they start another or more rounds. After each round, participants can view their game results in a summary chart displayed in their field of view or on their Apple Watch, showing their win rate, running distance, and calories burned.
@@
    Conclude Gameplay: Participants could play for their desired duration. However, as the MOFA games involved vigorous physical activity, players typically concluded gameplay after 3-5 minutes due to physical fatigue as a result of their body engagement. Additional bystanders sometimes wanted to join, and the researchers would rotate them in. This fluid grouping mimicked real-world street play, where people can spontaneously join or leave a game.
@@
    Contextual Inquiry: The researchers immediately asked if the participants had some more time for a brief interview. If not agreed, the participants leave the gameplay site; if agreed, one of the researchers interviews with the player or the player group. The other researcher continued to moderate the gameplay. In the interview, participants were invited to share their experiences and thoughts as contextual inquiry, one-on-one or in their original game groups of 2–3 people. While being brief, these interviews captured immediate reactions, which is valuable as participants could vividly recall their just-finished experience. We asked open-ended questions from a prepared interview question set demonstrated in Appendix A.
@@ 4.1.2. Data Collection
For the in-the-wild observations and post-play interviews, the following video recordings were collected.
[…]
Video recordings of AR spectator views.
For each session, a researcher holds an iPad with a spectator view, which shows the third-person perspective of player behaviors with augmented views of game content (see Figure  5). In total, we collected 265 minutes of recordings of players’ behavior data.
Interview recordings with players and spectators.
We collected a total number of 88 minutes of post-play interview video (see Table 2).
[…]
The contextual nature of these interviews allowed us to capture situated, embodied reflections in the moment, while the prepared questions list ensured consistency across sessions. Due to the fast-paced and public setting, especially in high-traffic environments like Expo (S2), Urban Parks (S3), and Public Transportation Buildings (S4), most participants only had a limited window for interaction, resulting in short interview durations (typically 2–4 minutes). Some participants who had consented to participate ultimately declined the interview due to time sensitivity or discomfort with extended conversation in public. We respected these preferences and included only the data of those who voluntarily continued. This pragmatic approach ensured ethical engagement while still allowing for the collection of rich, immediate feedback from diverse real-world contexts.
Over the study period, 65 individuals played the game across all sessions, of whom 32 volunteered for post-play interviews. These participants (labeled P1–P32 for those interviewed, and P33–P65 for additional players, who is not voluteered for post-play interviews) represented a broad demographic spectrum. The video data indicated a roughly balanced gender mix and a wide range of ages (teens with parents present, college students, adults, and a few seniors) and ethnic backgrounds.
4.1.3. Safety and Ethical Consideration
[…]
We don’t choose the dangerous place to play. Although there is no expectation of privacy when in public, we adhered to a number of principles during data collection. The locus of our capture was the play and we avoided recording people unnecessarily. Participants or bystanders could opt-out before interviews. No minors were interviewed; a few minors who tried the game did so under parental supervision, and we did not include their feedback in our dataset beyond general observations.
@@ 4.2. Desktop Research of Online Forum Reviews
In addition to on-site data, we wanted to capture broader public perceptions of IMRSP – including those of people who had not played the game but saw it online. To do this, we conducted a desktop research analysis of online discussions about our game probe. We posted a short gameplay video of MOFA and a description of the concept on a popular tech-enthusiast forum, inviting feedback and discussion. Over several weeks, this thread garnered considerable attention and commentary. We collected all responses from this forum thread, resulting in 14,919 words of raw text across numerous comments. We then filtered this dataset to focus on relevant content: specifically, comments that addressed the social implications, user experience, or concerns about MR street play (as opposed to off-topic remarks about technology or economics). After filtering out unrelated discussion, we had 1,325 words of substantive comments from 31 unique online users. These comments ranged from short quips to multi-sentence opinions (5 to 204 words each).
[…]
The online commenters had all viewed the same gameplay footage that our real-world participants experienced (via the posted video). Their perspectives are thus those of observers imagining the experience, rather than first-hand players. Not being physically present or involved with researchers, their feedback was uninfluenced by any demand characteristics or desire to please us. This gave us a more candid sense of excitement or skepticism that broader publics might have. Some commenters were enthusiastic about the idea, others were critical or concerned. We treated this as an additional qualitative dataset that complements the interview data: on-site participants provide insight into actual experience and immediate personal feelings, whereas online responses reveal more reflective or hypothetical viewpoints, including critiques from a social acceptance standpoint.
[…]
To understand broader societal perceptions, we analyzed online discussions featuring, referencing, or discussing our game. This approach provided perspectives from indirect observers. The analysis of online reviews helped us assess public receptivity, concerns, and attitudes towards IMRSP. Online reviews were collected from discussion threads on a tech-enthusiast forum (see Figure 6, totaling 14,919 words of comments. Game video links were provided in the initial post, ensuring all reviewers had seen the gameplay. After an initial review, 1,325 words of comments and discussions directly related to our game probes and research questions, focusing on social implications. We excluded discussions that are unrelated to mixed reality street play, such as general technology and economic topics. The relevant reviews ranged from 5 to 204 words in length and came from 31 unique users, labeled as L1-L31 (see Table 2). These online comments, unaffected by the researchers’ presence or developers’ interactions, offered unfiltered and naturalistic perspectives on the game and provided an additional layer of authentic feedback. The inclusion of online reviews complemented our research objective of understanding the holistic impact and acceptance of this novel gameplay format, incorporating both direct participant experiences and broader audience perspectives.
Figure 6. Comments on actual MOFA game video recording posted on an online tech-enthusiast forum
Table 2. Data collection from the online forum
@@ 4.3. Data Analysis
[…]
We conducted content analysis on the collected data. The video recordings were transcribed into texts for coding. We adopted an emergent coding technique, where we identify themes and patterns that arise naturally from the data rather than imposing predetermined codes. To begin with, three researchers familiarized themselves with the transcribed interview data and video recordings to gain a deep understanding of its content. Following this, one researcher completed the open coding on the data and the other two researchers reviewed the coded results, suggested edits, and reached an agreement on the coding results. Once initial codes were established, similar codes were grouped into broader themes. The researchers worked together to develop the themes and iteratively refined them to ensure that the themes accurately represent the core findings of the study. The themes were then validated by another researcher who was not involved in the coding process, cross-checking the themes against the data. All researchers agreed on the coding results and interpreted the findings based on the analysis, taking into account the occurrences of codes and themes as well as the contexts of the coded data. The results showed 121 unique codes that were organized into 4 themes: (1) social interactions, (2) social concerns, (3) gameplay experience, and (4) environment and location.
All interview video recordings were transcribed, and field notes from gameplay video observations were expanded into textual descriptions. For the online forum data, we already had the comments in text form. We analyzed the combined qualitative data using a content analysis with emergent thematic coding. We did not start with a rigid codebook; instead, three researchers independently read through subsets of the data (both interview transcripts, observation field notes, and online comments) to inductively identify recurring themes and patterns. After an initial open-coding pass, we compared notes and converged on a set of candidate themes capturing salient aspects of the experience. We then iteratively refined these, grouping related subthemes and ensuring they were grounded in the data.
[…]
During coding, we were careful to differentiate between on-site participant responses (P-codes) and online reviewer comments (L-codes). We created separate code categories when needed to reflect differences in perspective. For example, “awkwardness” was noted by both groups, but online commenters raised it more frequently and harshly; thus, we coded occurrences in the L data and P data and later compared them. We also quantified some occurrences to get a sense of prevalence (e.g. how many of the 63 total contributors mentioned a certain issue). These counts (noted as, say, 13/63) are used in the results to convey magnitude but are not treated as statistical measures.
[…]

### v2

@@ 3.3. System Implementation
To address the cost barrier of MRHMDs noted in Section 2.5, we employed HoloKit [^10] https://holokit.io/products/holokit-x, visited on October 25, 2024, an affordable headset (US\$129) (Hu et al., 2024c), which is a smartphone-based, open-source device providing a stereoscopic optical see-through mixed reality experience. This cost-effective solution facilitates multiplayer mixed reality experiences “in the wild”, enabling numerous passersby to engage with the gameplay.
The game probes were implemented using Unity 6 LTS [^11] https://unity.com/releases/unity-6, visited on October 25, 2024 and the HoloKit SDK [^12] https://github.com/holokit/holokit-unity-sdk, visited on October 25, 2024 based on ARKit [^13] https://developer.apple.com/augmented-reality/arkit/, visited on October 25, 2024. We deployed the game as an iOS application on iPhone 12 or newer models, which are inserted into the HoloKit. The headset displays stereoscopic renderings on the iPhone’s screen, allowing users to see stereoscopic 3D images overlaid on reality.
Interactions with augmented content were achieved through an Apple Watch. The watch’s motion sensors capture wrist movements for spell-casting, enabling ranged attacking gameplay without physical contact between players. This approach sidesteps the physical contact issue by allowing players to cast magic spells for intangible interaction with opponents. MOFA game probes feature various spell classes, each comprising a basic attack spell and an ultra spell. Players select one magic class per game round. While basic attack spells have minor variations in speed and size, ultra spells differ significantly, adding depth to the gameplay.
Figure 4. Spectators can use iPads to view the augmented reality experience from a third-person perspective and capture recordings of what they see.
We connect all co-located devices using the InstantCopresence toolkit (Hu et al., 2023a). The toolkit enables spontaneous mixed reality sessions when users scan a QR code on the host device. QR code scanning enables device co-location by calculating the relative position of client devices to the host’s QR code, allowing multiple devices to sync their coordinate systems based on their positions relative to the host origin. The network layer uses Multipeer Connectivity—Apple’s peer-to-peer wireless Wi-Fi and Bluetooth technology that powers AirDrop—to provide local networking with latency under 20 ms. This enables street play anywhere, even in rural areas without cellular or Wi-Fi coverage. An auto-sync system keeps virtual object data consistent across devices, ensuring all participants see the same virtual elements. Spectators can view the augmented content from a third-person perspective using mobile devices like iPads (see Fig. 4).
@@ 4. Method
To investigate IMRSP, we conducted an in-the-wild study (Benford et al., 2013) combining direct gameplay video observations with immediate post-play contextual inquiries (Duda et al., 2020). To gain a broader perspective, we shared in-the-wild gameplay recordings on a tech‑enthusiast forum. Following this, we conducted desktop research to gather and analyze online reviews and comments about the experience. This combined approach allowed us to capture both firsthand player experiences and wider public perceptions. Our study was approved by the authors’ institutional ethics review, and all participants gave informed consent prior to participation.
@@ 4.1. In-the-Wild Field Study
We conducted a series of in-the-wild play sessions over three years, in which members of the public were invited to try MOFA. These sessions occurred as pop-up events at 20 different locations across North America, Asia, Europe, and Oceania, spanning settings such as university campuses, tech expos, urban parks, public building courtyards, and natural wilderness sites.
4.1.1. Scenarios
For this paper, we focus on five representative scenarios (see Fig. 5): public areas of university campus (S1), exhibition spaces at expos (S2), urban parks (S3), public building atria (S4), and scenic mountainside in natural wilderness (S5). These diverse contexts allowed us to examine IMRSP across varying degrees of publicness, crowd density, and thematic fit.
S1 Campus: Semi-public university courtyards have moderate foot traffic (5-10 passersby per session). The steady trickle of students and faculty creates a sociable backdrop for exploration. Studying IMRSP here reveals how ambient social presence affects curiosity and engagement with the game in semi-public familiar settings.
S2 Expos: Expo and conference halls present a high-visibility, demonstration‑friendly environment characterized by dense crowds (30-50 observers per session), rapid turnover of visitors, and visually busy booth layouts. IMRSP deployment here uncovers strategies for maintaining engagement amid rapid attendee turnover, where audiences are generally more curious and have a higher threshold for social acceptance.
S3 Urban Parks: Public outdoor areas offer open space, greenery, and diverse pedestrian flow (10-20 bystanders per session), including Dumbo Park in Brooklyn and Washington Square Park in Manhattan. Testing in these environments shows how natural elements and unstructured layouts affect spontaneous participation in leisure settings.
S4 Public Buildings: This scenario is characterized by iconic indoor spaces with continuous pedestrian flows, echoic soundscapes, and high visibility (20-50 people per session), such as Oculus Transportation Hub and Grand Central Terminal. Gameplay in these settings highlights necessary design adaptations for embodied interactions under public social pressure.
S5 Natural Wilderness: This is a typical scenario that has unbounded landscapes with open ground, vertical space, and minimal bystander presence (0-2 bystanders per session), such as the foot of mountains in national parks. Exploring IMRSP here reveals how the experience transforms when social constraints disappear and nature becomes the primary contextual element.
Figure 5. Five representative scenarios of volunteered participation and interviews. S1: Campus; S2: Expos; S3: Urban Parks; S4: Public Buildings; S5: Natural Wilderness.
Figure 6. Data collection from the in-the-wild study.
@@ 4.1.2. Procedure
In order to collect ecologically valid MR street play data, we used an opportunistic recruitment strategy, following the method used in previous works (Denning et al., 2014; O’Hagan et al., 2021). This recruitment approach is common in in-the-wild research (Rogers and Marshall, 2017), which yielded relatively short engagements due to the nature of interrupting people during their normal routine. Nevertheless, it encompassed a broad spectrum of individuals who encountered this MR street play in those public spaces, thereby enhancing the diversity of our sample. The procedure for a typical street‑play session is described below.
@@
    Configure the play area: In each pop-up or in-booth show, two or three researchers ran the session in a space of at least 3 × 3 meters in a public area, ensuring minimal disruption to pedestrian flow. A “Video Recording In Progress” sign was displayed. For Expo scenarios (S2), we demonstrated the game at the booth and via pop‑ups in lobbies or near entrances during breaks, following our in‑the‑wild deployment approach.
@@
    Engage passersby: Pedestrians were drawn to the ongoing gameplay, introduction videos, and spectator footage displayed on iPads and portable screens. As passersby were drawn to the play and began to interact with the researchers, we invited them to try the game.
@@
    Obtain informed consent: Interested participants signed consent forms on an iPad or paper.
@@
    Equip participants for a game mode: To facilitate a smooth and context-aware play experience, researchers assisted participants with initial setup, including pairing players’ devices, wearing the MR headset, securing the motion-tracking watch, and selecting their preferred virtual spells. The choice of MOFA game mode was guided by a combination of player preferences, social context, site constraints, and time availability (see Table 1). Participants were introduced to the available game options, and the most suitable mode was recommended based on the situation. The adaptive approach maintained ecological validity while ensuring that each participant could meaningfully engage within the constraints of the public and often time-limited settings.
@@
    • When another participant became available, they were often invited to join in “The Duel”, which requires two players to engage in competitive interactions. Participants who arrived together—such as friends or colleagues—were directly introduced to “The Duel” to leverage their existing social rapport.
    • When space permitted extended movement (e.g., campus playgrounds or natural wilderness) and participants had the time and physical ability, researchers suggested “The Dragon”, a longer (about 5 minutes) and more physically demanding cooperative game mode.
@@
    Facilitate gameplay sessions: The first 90 seconds (the first round) served as a tutorial. Once participants learned to attack and dodge in MR, they played one or more additional rounds. After each round, participants could view results in a summary chart in MR view or on Apple Watch, including score, hit rate, distance moved, and calories burned.
@@
    Conclude gameplay: Participants could play for their desired duration. However, as the MOFA games involved vigorous physical activity, players typically concluded after 3–5 minutes due to physical fatigue from bodily exertion. When bystanders wished to join, researchers rotated them into the session. This fluid grouping mimicked real-world street play, where people can spontaneously join or leave a game.
@@
    Contextual inquiry: Researchers immediately asked participants if they had time for a brief interview. If they declined, the researchers let them leave. Otherwise, one researcher interviewed either individuals or player groups while another moderated gameplay. During the interview, participants shared their experiences and thoughts either one-on-one or in their original game groups of 2–3 people. Although brief, these interviews captured valuable immediate reactions while participants could still vividly recall their fresh experience. We used a prepared set of open‑ended questions listed in Appendix A.
@@ 4.1.3. Data Collection
For the in-the-wild observations and post-play interviews, two types of video recordings were collected: AR spectator views and interview recordings. For each session, a researcher held an iPad with a spectator view, which shows the third-person perspective of player behaviors with augmented views of game content (see Fig. 4). In total, we collected 265 minutes of recordings of players’ behavior data. In addition, we collected a total number of 88 minutes of post-play interview video, as detailed in Fig. 6.
The contextual nature of these interviews allowed us to capture situated, embodied reflections in the moment, while the prepared questions list ensured consistency across sessions. Due to the fast-paced and public setting, especially in high-traffic environments like Expo (S2), Urban Parks (S3), and Public Buildings (S4), most participants only had a limited window for interaction, resulting in short interview durations (typically 2–4 minutes). Some participants who had consented to participate ultimately declined the interview due to time sensitivity or discomfort with extended conversation in public. We respected these preferences and included only the data of those who voluntarily continued. This pragmatic approach ensured ethical engagement while still allowing for the collection of rich, immediate feedback from diverse real-world contexts.
Over the study period, 65 individuals played the game across all sessions, of whom 32 volunteered for post-play interviews. These participants (labeled P1–P32 for those interviewed, and P33–P65 for additional players, who did not participate in the post-play interviews) represented a broad demographic spectrum. The video data indicated a roughly balanced gender mix and a wide range of ages (teens with parents present, college students, adults, and a few seniors) and ethnic backgrounds.
4.1.4. Safety and Ethical Considerations
We avoided any hazardous locations. Although privacy is not expected in public spaces, we followed strict ethical guidelines during data collection. The focus of our capture was the play, and we avoided recording people unnecessarily. Participants or bystanders could opt out before interviews. No minors were interviewed; a few minors who tried the game did so under parental supervision, but we did not include their feedback in our dataset beyond general observations.
@@ 4.2. Desktop Research of Online Forum Reviews
In addition to on-site data, we wanted to capture broader public perceptions of IMRSP—including those of people who had not played the game but saw it online. To do this, we conducted a desktop research analysis of online discussions about our game probe. We posted a short gameplay video of MOFA and a description of the concept on a popular tech-enthusiast forum, inviting feedback and discussion. Over several weeks, this thread garnered considerable attention and commentary. We collected all responses from this forum thread, resulting in 14,919 words of raw text across numerous comments. Fig. 7 shows some snippets of the online forum threads. We then filtered this dataset to focus on relevant content, specifically, comments that addressed the social implications, user experience, or concerns about MR street play (as opposed to off-topic remarks about technology or economics). After filtering out unrelated discussion, we had 1,325 words of substantive comments from 31 unique online respondents (L1–L31). These comments ranged from short quips to multi-sentence opinions (5 to 204 words each).
The online respondents had all viewed the same gameplay footage that our real-world participants had experienced (via the posted video). Their perspectives are thus those of observers envisioning the experience, rather than firsthand players. Not being physically present or involved with researchers, their feedback was not influenced by any demand characteristics or desire to please us. This provided us with more candid feedback and skepticism than we might have received from the public onsite. Some commenters were enthusiastic about the idea, while others were critical or concerned. We treated this as an additional qualitative dataset that complements the interview data: on-site participants provide insight into actual experience and immediate personal feelings, whereas online responses reveal more reflective or hypothetical viewpoints, including critiques from a social acceptance standpoint.
Figure 7. Comments on actual MOFA game video recording posted on an online tech-enthusiast forum.
@@ 4.3. Data Analysis
All interview video recordings were transcribed, and field notes from gameplay video observations were expanded into textual descriptions. For the online forum data, we already had the comments in text form. We analyzed the combined qualitative data using a content analysis with emergent thematic coding. We did not start with a rigid codebook; instead, three researchers independently read through subsets of the data (including transcripts of the observation and interview videos and online comments) to inductively identify recurring themes and patterns. After an initial open-coding pass, we compared notes and converged on a set of candidate themes capturing salient aspects of the experience. We then iteratively refined these, grouping related subthemes and ensuring they were grounded in the data.
During coding, we were careful to differentiate between on-site participant responses (P-codes) and online reviewer comments (L-codes). We created separate code categories when needed to reflect differences in perspective.
Throughout the analysis, we practiced researcher triangulation: multiple coders cross-checked each other’s assigned codes and interpretations. Differences were discussed and resolved. We also revisited raw videos to verify behaviors for certain themes. This process ensured that our findings are solidly grounded in the empirical evidence. We group our findings by four broad thematic categories: (1) Social Interactions, (2) Social Concerns, (3) Gameplay Experience, and (4) Environment and Location.